C=500+0.5YD
I=100
T=80
G=200
Solve the good market equilibrium
1
Expert's answer
2019-08-22T14:03:44-0400

Total demand: Z = C + I + G

Good market equilibrium: Y = Z

So, Y = C + I + G 

Y(D) = Y – T 

Y = 500 + (0.5) (Y – 80) + 100 + 200

Y = (0.5) Y + 760 Y = 1520 (equilibrium output) 

YD = 1520 – 80 = 1440 (disposable income)

C = 500 + (0.5) (1440) C = 1220 (equilibrium consumption)

Z = 1220 + 100 + 200 Z = 1520 (total demand)